OpenStudy (drewbooks):

What are the real roots of the given equation? 3(x-4)^(4/3) + 16 = 64

OpenStudy (anonymous):

x=12

OpenStudy (drewbooks):

Thank you. How did you solve that?

OpenStudy (anonymous):

mathpapa.com

OpenStudy (drewbooks):

Oh. Does someone else have an explanation?

Nnesha (nnesha):

math papa

Nnesha (nnesha):

yes ofc drew we will explain

OpenStudy (johnweldon1993):

\[\large 3(x - 4)^{4/3} + 16 = 64\] Subtract 16 from both sides of the equation \[\large 3(x-4)^{4/3} = 48\] Divide both sides by 3 \[\large (x - 4)^{4/3} = 16\] Can you take it from there

OpenStudy (drewbooks):

Got it. I hadn't seen a question on finding "real roots" before, so I didn't know what to do. Thank you for the explanation.

OpenStudy (anonymous):

make sure to use PEMDAS

OpenStudy (johnweldon1993):

Right, that is just their way of saying "what value(s) of 'x' will make this a correct statement"

OpenStudy (drewbooks):

Cheers.

Nnesha (nnesha):

@TJEscobar i like that website math "papa" :P ahhah nice name :-)

OpenStudy (anonymous):

@Nnesha me Too!! I use it whenever I get stuck!! I'm actually new here so Iv'e been figuring stuff out by myself up until 15 minutes ago!! XDD

Nnesha (nnesha):

yeah but this is openSTUDY we must hve to explain :-)

OpenStudy (anonymous):

true